OpenStudy (anonymous):

Three successive lines in the atomic spectrum of hydrogen have wave lenghts 18746A , 12815A and 10935A calculate the series limit for the spectral series to which the above lines belong?

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

@phi @mukushla @Callisto

OpenStudy (ghazi):

i think you've to use Rydberg's formula

OpenStudy (ghazi):

\[\frac{ 1 }{ \lambda }= R_{h}*(\frac { 1 }{ n _{1}^2 }-\frac{ 1 }{ n _{2}^2 })\]

OpenStudy (dls):

8208 Amperes?

OpenStudy (ghazi):

??

OpenStudy (dls):

what's the answer?

OpenStudy (dls):

@Yahoo!

OpenStudy (anonymous):

@DLS u r correct

OpenStudy (anonymous):

but it is no ampere it is Amstrong

OpenStudy (dls):

ah sorry yeah armstrong

OpenStudy (anonymous):

Can u show hw u got it

OpenStudy (dls):

no :P

OpenStudy (anonymous):

why

OpenStudy (dls):

XD wher edid you try?

OpenStudy (anonymous):

i have No idea..))

OpenStudy (anonymous):

i nw Rydberg's formula....)) but i dont find a way to use it here

OpenStudy (anonymous):

@ghazi any idea

OpenStudy (anonymous):

hm....Can u show...@DLS

OpenStudy (dls):

See you are given 3 successive wavelengths Apply that formula :)

OpenStudy (anonymous):

and we have to find n..))

OpenStudy (dls):

you wil lget 3 equations

OpenStudy (anonymous):

@ghazi the answer key tells like this 912* 9 = 8208 A

OpenStudy (anonymous):

and wat does they really mean...

OpenStudy (ghazi):

i think it's wavelength multiplied by 3^2 that is value of n..but i ain't sure

OpenStudy (anonymous):

and wat is 912??

OpenStudy (ghazi):

\[\frac{ 1 }{ 18746*10^{-10} }= R _{h}(\frac{ 1 }{ n _{1}^2 }-\frac{ 1 }{ n _{2}^2 })\] you'll find three equations by this and solve for it

OpenStudy (anonymous):

lol....ok...912..))) rydberg consatnt ---- 109677

OpenStudy (dls):

like i told you :)

OpenStudy (ghazi):

yess...Rh= 1.0973*10^7 m

OpenStudy (anonymous):

and wat is 912

OpenStudy (dls):

a term :)